Sam is 15 years younger than Jane. 3 years ago, Jane's age was 2 times Sam's age. How old is Sam now?
Answer by JulietG(1812)

S = J-15
2(S-3) = J-3
Distribute
2S - 6 = J - 3
Substitute the known value of S from the top equation into the bottom
2S-6 = J-3
2(J-15)-6 = J-3
Distribute
2J-30-6 = J-3
2J-36 = J-3
Subtract J from each side
J-36 = -3
Add 36 to each side
J = 33
If Jane is 33, then same is 18.

Prove it: 3 years ago, Jane's age was 2 times Sam's age.
3 years ago, Jane was 33-3, or 30
3 years ago, Sam was 18-3, or 15
30 is twice 15.